SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (August 22, 2006)

Attendees: Rainer Bartoldus, Al Eisner, Arik Kreisel, Selina Li, Rahmat Rahmat, Dong Su, Mark Tibbetts

Operations

  • BaBar/PEP shutdown began earlier than scheduled due to a fire around IR4 region last Thursday swing shift. The last physics run in run 5b is 67707.
  • The TSF/BLT/ZPD crates were turned off yesterday because of the cooling work going on today. They will be remained off for the next one to two weeks or till further notice since we will be taking some of the TSF boards to our teststand for firmware testing.

EMT

  • Mark will be flying back to UK next week and be back to SLAC again on Sept.24th to take over Dan's responsibilities.

DCT

  • Marc pointed out that we need to track down to which and how many TSF's are out of sync with the rest of them, and in what way. We could probe the test points on the TSFi/ZPDi to see what the framebit pulses are like, and if they jump around compared to each other. Su Dong suggested that we can study the collision data that we took with the version of the firmware which caused the ZPD damage. We could look at the TSF -> ZPD transmission. We will take the boards that showed damage to the teststand (rather than IR2) for further investigation.

Online

  • Rainer found a bug fixed (to prevent memory leak) for ZPD configuration that didn't go in the release.

L3

  • Arik reported a minor improvement for the seed list in L3 tracking. He also generated some BBbar samples in Moose to study the impact on tracking efficiency.
  • Rainer had an idea about relaxing the Pt cut or perhaps putting the Pt cut in the configuration database since it would be cleaner but more work required since new object is required.

Data Quality and Monitoring

  • Rahmat showed the lumi ratio plots and strip charts. It was suggested to use run number or time/date on the x-axis for the lumi ratio strip charts. Rainer will spend some time to help Rahmat with this.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.